OpenStudy (anonymous):

CUBE ROOTS OF -8: I find that r =sqrt(-8^2 +0)= 8. And Arg IzI = pi or 180 degrees. Now, when k=0, I am getting 2[ cos0 + iSin0), when writing in complex form this will give me 2, cos 0=1! or am I missing something, because, 2 x 2 x 2 =8 not -8?

ganeshie8 (ganeshie8):

\(\large - 8 = 8(\cos \pi + i \sin \pi) = 8e^{i \pi }\)

ganeshie8 (ganeshie8):

cube roots of -8 = \(\large \sqrt[3]{8} ~e^{i \frac{\pi + 2k \pi}{3} } = 2 ~e^{i \frac{\pi + 2k \pi}{3} } = 2 \left[ \cos ( \frac{\pi + 2k\pi}{3}) + i \sin ( \frac{\pi + 2k\pi}{3}) \right] \) \(0 \le k \lt 3 \)

OpenStudy (anonymous):

so when k =0?

OpenStudy (anonymous):

that is the part that i don't understand

OpenStudy (anonymous):

the result i get in complex form is 2(1 +0i)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large 2 \left[ \cos ( \frac{\pi + 2k\pi}{3}) + i \sin ( \frac{\pi + 2k\pi}{3}) \right]\] plugin \(k = 0\) \[\large 2 \left[ \cos ( \frac{\pi }{3}) + i \sin ( \frac{\pi}{3}) \right]\]

ganeshie8 (ganeshie8):

thats another cube root for -8, cube it and see if u get -8